OpenStudy (anonymous):

7. A ball dropped from a certain height rises after bouncing to 25 % of its original height. After bouncing if the ball rises to 6 ft. What is the height from which it was dropped?

OpenStudy (alfie):

6ft : 25% = x ft : 100% 600/25 = 24ft

OpenStudy (anonymous):

ok--i'm trying to figure out out how to determine which is the base, rate and amount

OpenStudy (alfie):

Well, I guess you are trying to do a kinematic's exercise now. Since we are talking of a ball launched vertically, you just have one dimension 24ft = your initial point -9,81m/s^2 = acceleration 0 = your final point You can find pretty much all you need now ;).. good luck

OpenStudy (anonymous):

h= original height so 25%h=0.25h=6 feet rises h=6/0.25=24 feet

OpenStudy (anonymous):

if its in feet g=32.2 ft/sec^2

OpenStudy (anonymous):

good luck bill